### v1

# ViewMask-1-to-3: Multi-View Consistent Image Generation via Multimodal Diffusion Models

Abstract
Multi-view image generation from a single image and text description remains challenging due to the difficulty of maintaining geometric consistency across different viewpoints. Existing approaches typically rely on 3D-aware architectures or specialized diffusion models that require extensive multi-view training data and complex geometric priors. In this work, we introduce ViewMask-1-to-3, a pioneering approach to apply discrete diffusion models to multi-view image generation. Unlike continuous diffusion methods that operate in latent spaces, ViewMask-1-to-3 formulates multi-view synthesis as a discrete sequence modeling problem, where each viewpoint is represented as visual tokens obtained through MAGVIT-v2 tokenization. By unifying language and vision through masked token prediction, our approach enables progressive generation of multiple viewpoints through iterative token unmasking with text input. ViewMask-1-to-3 achieves cross-view consistency through simple random masking combined with self-attention, eliminating the requirement for complex 3D geometric constraints or specialized attention architectures. Our approach demonstrates that discrete diffusion provides a viable and simple alternative to existing multi-view generation methods, ranking first on average across GSO and 3D-FUTURE datasets in terms of PSNR, SSIM, and LPIPS, while maintaining architectural simplicity.
Figure 1: Comparison of multi-view image generation approaches. (Left-Top) 3D representation methods require expensive 3D reconstruction and rendering. (Left-Bottom) Camera-conditioned diffusion methods need precise camera parameters. (Left-middle) Image editing models generate views sequentially without consistency guarantees. (Right) Our discrete diffusion approach enables parallel multi-view generation through unified sequence modeling with flexible text-based control.
1 Introduction
Multi-view image generation has emerged as a fundamental challenge [he2024head360, liu2024mvsgaussian, he2024emotalk3d], with applications spanning virtual reality [zhou2025cross] and 3D reconstruction [alzayer2024seeing]. The task aims to generate multiple consistent viewpoints of an object or scene from limited input—typically a single image and accompanying text description. Computational approaches face significant challenges in maintaining geometric consistency, preserving fine-grained details, and ensuring semantic coherence across viewpoints.
Traditional multi-view synthesis approaches primarily rely on explicit 3D representations or geometry-aware neural architectures. 3D-aware GANs [chan2021pi, schwarz2020graf] leverage neural radiance fields or voxel grids to enforce geometric constraints, while recent diffusion-based methods [shi2024mvdream, liu2024syncdreamer] employ specialized architectures with view-dependent conditioning. However, these approaches often require extensive multi-view training datasets, complex geometric priors, or computationally expensive 3D intermediate representations. Moreover, they typically operate in continuous latent spaces, limiting precise controllability over the generation process.
Recent advances in discrete diffusion models [nie2025large, ye2025dream, sahoo2024simple, you2025lladav] have demonstrated remarkable success in language modeling and multimodal understanding by treating generation as progressive token unmasking, models like LLaDA [nie2025large] for language generation, LLaDA-V [you2025lladav] for vision-language understanding, MMaDA [yang2025mmada] and Lumina-DiMOO [xin2025luminadimooomnidiffusionlarge] for multimodal generation, enable parallel decoding and bidirectional context modeling. These models operate by progressively revealing masked tokens in a learned sequence, enabling iterative refinement and controllable generation.
Discrete diffusion models offer two significant benefits: (1) Unified architecture across modalities, aligning with the industry shift toward integrated frameworks and enabling seamless cross-modal integration [huunified2023udd, yang2025mmada]. (2) Efficient parallel processing that surpasses sequential autoregressive methods in both speed and flexibility through simultaneous token prediction. The inherent bidirectional attention naturally captures contextual relationships, making discrete diffusion uniquely suited for tasks requiring consistency across complex dependencies [chang2022maskgit, xin2025luminadimooomnidiffusionlarge].
In this work, we introduce ViewMask-1-to-3, a novel approach that adapts discrete diffusion models to multi-view image generation. Our key insight is that multi-view generation can be naturally formulated as a sequence modeling problem, where each viewpoint is represented as a sequence of discrete visual tokens obtained through MAGVIT-v2 [yu2024language] tokenization. By extending masked token prediction from language to the vision, ViewMask-1-to-3 enables progressive generation of multiple viewpoints through iterative token unmasking.
Unlike existing approaches [liu2023zero, kwak2023vivid, shi2023zero123++, zhang2025ar] that rely on precise camera parameters [liu2023zero, zhang2025ar], 3D reconstruction rendering [tangdreamgaussian, liu2024syncdreamer], or large image editing models that can only generate a single view at once with coarse view control, ViewMask-1-to-3 achieves cross-view consistency through simple random masking combined with bidirectional attention, as illustrated in Figure 1. This approach allows for parallel refinement of visual tokens across all viewpoints, while the iterative nature of masked diffusion provides opportunities for progressive quality improvement during the generation process. Our contributions are as follows:
  • We present a novel discrete diffusion method for multi-view image generation task, demonstrating how masked token prediction can effectively address visual synthesis across multiple viewpoints.
  • The proposed simple masking strategy, when combined with bidirectional attention mechanisms, naturally achieves cross-view consistency without requiring complex 3D geometric priors or specialized architectures.
  • We evaluate ViewMask-1-to-3 on standard multi-view generation benchmarks, achieving the best average performance on both GSO and 3D-FUTURE datasets, establishing discrete diffusion as a viable alternative to existing continuous diffusion and 3D-aware methods.
2 Related Work
2.1 Multi-View Image Generation
3D Representation-Based Methods
Early approaches ensure geometric consistency through explicit 3D modeling. 3D-aware GANs such as $\pi$-GAN [chan2021pi] and EG3D [chan2022efficient] leverage neural radiance fields or tri-plane representations for view-consistent synthesis. More recently, DreamGaussian [tangdreamgaussian] employs 3D Gaussian Splatting for efficient content creation; TRELLIS [xiang2024structured] adopts flow matching to obtain structured 3D latents and then extracts multi-view images; MVDream [shi2024mvdream] and SyncDreamer [liu2024syncdreamer] integrate NeRF [mildenhall2021nerf] into diffusion frameworks with synchronized denoising across viewpoints. While achieving strong geometric coherence, these methods require computationally expensive 3D representations and specialized rendering pipelines that increase system complexity.
Camera-Conditioned Diffusion Methods
An alternative approach generates multi-view images directly in 2D space through camera-conditioned diffusion. Zero-1-to-3 family [liu2023zero, deitke2023objaverse, shi2023zero123++] pioneered fine-tuning pretrained diffusion models with camera pose conditioning for zero-shot novel view synthesis. Recent extensions include ViVid-1-to-3 [kwak2023vivid], which leverages video diffusion for temporal coherence; AR-1-to-3 [zhang2025ar], which introduces autoregressive viewpoint generation; EpiDiff [huang2024epidiffenhancingmultiviewsynthesis], which enforces epipolar constraints through specialized attention mechanisms. However, these methods rely on explicit camera parameterization and specialized pose conditioning architectures, requiring precise viewpoint specifications and limiting flexibility in view specification through natural language.
Image Editing Methods
Recent methods attempt to generate multi-view images via image editing strategies. These approaches leverage various techniques: MV-Adapter [huang2024mvadapter] transforms text-to-image models into multi-view generators without altering their structure; ImageBrush [yang2023imagebrush] employs visual instructions rather than text for precise manipulation; SceneScape [fridman2023scenescape] ensures geometric consistency through depth-aware generation; and Lumina-DiMOO [xin2025luminadimooomnidiffusionlarge] introduces discrete diffusion in image editing with impressive results. While these methods rely on specialized editing strategies, ViewMask-1-to-3 naturally achieves cross-view consistency through token-level modeling using discrete diffusion, offering a more unified and elegant solution to multi-view synthesis.
2.2 Discrete Diffusion Models
D3PM [austin2021structured] established the foundational framework for discrete diffusion, extending traditional diffusion to discrete state spaces through structured denoising processes. DiffusionBERT [he2023diffusionbert] demonstrated that pre-trained BERT models could be adapted for text generation through discrete diffusion training. A significant breakthrough came with LLaDA [nie2025large], which showed that masked diffusion models can match autoregressive models like LLaMA3-8B in language generation while offering parallel decoding and bidirectional context modeling. LLaDA-V [you2025lladav] extended this framework to multimodal understanding, demonstrating the potential of discrete diffusion in vision-language tasks. These advances have established discrete diffusion as a viable alternative to autoregressive approaches, for tasks requiring controllable and parallel generation.
2.3 Visual Tokenization
The effectiveness of discrete diffusion for visual tasks critically depends on visual tokenization quality. VQVAE [oord2017neural] introduced discrete visual representations through vector quantization, laying the groundwork for discrete visual modeling. MaskGIT [chang2022maskgit] applied masked token prediction to image generation using vector-quantized representations. MAGVIT-v2 [yu2024language] enabling high-quality discrete tokenization for both images and videos through lookup-free quantization and architectural improvements. The success of these tokenizers has made discrete approaches increasingly competitive with continuous methods in visual generation tasks.
Figure 2: Overview of our three-stage training framework. (I) Stage 1: Pretraining aligns text and image modalities using masked image-caption sequences. (II) Stage 2: Image-to-Multi-View (I2MV) masks the target view and learns to reconstruct it conditioned on a reference view, thereby acquiring multi-view generation capability. (III) Stage 3: Text-to-Multi-View (T2MV) masks all views and learns to generate multiple views of an object conditioned on a text description.
3 ViewMask-1-to-3
3.1 Overview
Given a single input image $I_{0}\in\mathbb{R}^{H\times W\times 3}$ and an optional text description $T$, our goal is to generate $N=3$ consistent target views $\{I_{1},I_{2},I_{3}\}$ representing the same object from different viewpoints. Beyond image-conditioned view synthesis, our model also achieve text-only generation: provided solely with a description $T$, it can synthesize $N=4$ new views $\{I_{1},I_{2},I_{3},I_{4}\}$ that correspond to distinct viewpoints, even in the absence of an input image. This unified formulation enables both image-to-multi-view and text-to-multi-view generation within a single framework. Unlike Zero-1-to-3 that operates in continuous latent spaces, we formulate multi-view generation as a discrete sequence modeling problem. By representing each view as a sequence of discrete visual tokens, we can leverage the parallel processing and bidirectional context modeling capabilities of masked diffusion models.
3.2 Model Design
Our approach consists of three main components: (1) Visual Tokenization, where we convert images to discrete token sequences using MAGVIT-v2; (2) Cross-View Sequence Construction, where we build unified sequences encoding multiple viewpoints with special tokens; and (3) Masked Diffusion Training, where we learn to predict masked visual tokens through iterative refinement.
Visual Tokenization
We adopt MAGVIT-v2 [yu2024language] as our visual tokenizer due to its superior performance in discrete visual representation. For an input image $I_{i}$, the tokenizer produces a sequence of discrete visual tokens,
$V^{(i)}=\text{Tokenize}(I_{i})=\{v^{(i)}_{1},v^{(i)}_{2},\ldots,v^{(i)}_{L}\}$, (1)
where $V^{(i)}\in\mathcal{V}^{L}$ represents the token sequence, $\mathcal{V}$ is the visual vocabulary with $|\mathcal{V}|=2^{18}$ tokens, and $L$ is the sequence length. Each token $v^{(i)}_{j}$ corresponds to a spatial region of the image and encodes local visual features. The discrete representation enables us to apply language modeling techniques directly to visual content while maintaining reconstruction quality through the learned tokenizer.
The tokenized sequences are then embedded into the language model’s embedding space:
$\mathbf{e}^{(i)}=\text{Embed}(V^{(i)})\in\mathbb{R}^{L\times d},$ (2)
where $d$ is the embedding dimension of the transformer.
Cross-View Sequence Construction
The unified sequence construction is a core step of ViewMask-1-to-3, which enables coordinated generation across multiple viewpoints under both image-to-multi-view (I2MV) and text-to-multi-view (T2MV) settings. By using the same masked-sequence format, we minimize the structural gap between I2MV and T2MV inputs, enabling the model to treat both tasks similarly (see Section 3.3 for details).
Cross-View Masking Strategy
During training, we employ a simple yet effective random masking strategy. For each training sample, target tokens are first selected based on the current training stage (see Section 3.3), and a masking ratio $r\sim\text{Uniform}(0,1)$ is sampled to randomly replace a subset of them with a special [MASK] token.
Figure 3: Overview of inference framework. (a) I2MV: The tokenized prompt and reference view provide a template for generating three consistent output images from different viewpoints. (b) T2MV: Tokenized prompt and description, providing a template for generating four output images. At each timestep t, the tokens are further predicted, and under the guidance of the mask schedule, remasking is performed based on confidence.
3.3 Multi-Stage Training Strategy
As shown in Fig. 2, our training framework comprises three stages.
Stage1: Pretraing for Multimodal Alignment
To enable the LLADA model to learn representations of image tokens, we leverage image-caption pairs to align the two modalities. The input sequence is formatted as:
[…]
where $\text{T2I}]$ indicates the image-caption alignment, $\text{SOT}]$ and $\text{EOT}]$ denote the start and end of the text, and $\text{SOI}]$ and $\text{EOI}]$ denote the start and end of the image. Both caption tokens and image tokens are randomly masked with a certain probability during training. This design encourages the model to fully capture the co-occurring semantic information in image–caption pairs, thereby aligning the representations of the two modalities.
We construct a multimodal image–caption dataset by combining ImageNet [deng2009imagenet] images paired with captions from ImageNet-1K-VL-Enriched [VisualLayer2024ImageNetEnriched] and the LLaVA Visual Instruct CC3M Pretrain dataset [liu2023llava], which is a subset released by LLaVA from the CC-3M corpus, to facilitate cross-modal alignment.
Stage2: Image to Multi-View
To enable the LLADA model to learn representations of image tokens, we leverage image–caption pairs to align the two modalities. The input sequence is formatted as:
$\text{I2MV}]~[\text{SOT}]~\textit{P.}~[\text{EOT}]~[\text{SOI}]~\textit{R.}~[\text{EOI}]~[\text{SOI}]{\color[rgb]{0.62890625,0,0}~\textit{G}_{1}~\textit{G}_{2}~\textit{G}_{3}}~[\text{EOI}],$
where $\text{I2MV}]$ indicates the image-to-multi-view task, P. represents the prompt tokens, R. represents the reference image tokens, and $\textit{G}_{i}$ represents the generated image tokens. For brevity, the repeated $\text{EOI}]~[\text{SOI}]$ between consecutive $\textit{G}_{i}$ are omitted. $\textit{G}_{i}$ are randomly masked with a certain probability during training.
At this stage, the model is trained on a curated subset of Objaverse [deitke2023objaverse] and Habitat Synthetic Scene Dataset (HSSD) [khanna2023hssd]. Following the data split provided by Objaverse++ [lin2025objaversecurated3dobject], we removed samples labeled as Low Quality (No semantic meaning. Objects that annotators cannot identify or are corrupted), to ensure more stable training.
Stage3: Text to Multi-View
To extend the I2MV paradigm to T2MV, we incorporate an additional textual description specifying the target views immediately after the prompt. Accordingly, the input sequence is formatted as:
$\text{T2MV}]~[\text{SOT}]~\textit{P.}~\textit{D.}~[\text{EOT}]~[\text{SOI}]{\color[rgb]{0.62890625,0,0}~\textit{G}_{1}~\textit{G}_{2}~\textit{G}_{3}~\textit{G}_{4}}~[\text{EOI}],$
where $\text{T2MV}]$ indicates the text-to-multi-view task, D. represents the description tokens. For brevity, the repeated $\text{EOI}]~[\text{SOI}]$ between consecutive $\textit{G}_{i}$ are omitted. $\textit{G}_{i}$ are randomly masked with a certain probability during training.
We enhance Objaverse with Cap3D [luo2023scalable, luo2024view], which offers detailed textual descriptions for 3D assets.
Loss
ViewMask-1-to-3 is trained using a cross-entropy loss that predicts original tokens at masked positions.
$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{3}\sum_{j\in\mathcal{M}_{i}}\log P(v^{(i)}_{j}|s_{\setminus\mathcal{M}}),$ (3)
where $\mathcal{M}_{i}$ denotes the set of masked positions in view $i$, and $s_{\setminus\mathcal{M}}$ represents the sequence with masked tokens. This objective encourages the model to reconstruct original visual tokens while leveraging unmasked context from the input image, text description, and partially observed target views. The cross-attention mechanism naturally enables information flow between different modalities and viewpoints.
3.4 Inference Process
Construction of Input Sequences
During inference, we perform iterative denoising starting from fully masked target views, as illustrated in Fig. 3. For I2MV, the model is provided with a tokenized prompt and the reference view. The template is designed to generate three output images. For T2MV, the input consists of a tokenized prompt and a textual description of the desired view. This template supports the generation of four output images.
Predict and Remask
We employ an iterative denoising-based sampling strategy. The model progressively predicts token distributions over $T$ iterations. At each step, the model performs a forward pass to obtain logits for all masked positions and estimates the confidence of each token prediction. A scheduling function (cosine, linear, or quadratic) determines the number of tokens to be re-masked in the next step, where low-confidence tokens are re-masked while high-confidence ones are retained. This predict–re-mask loop continues until all masked tokens are generated, progressively refining the sequence toward the final output. We further discuss the effects of different scheduling strategies in the ablation study (see Section 4.3).
4 Experiments
Figure 4: Qualitative comparison on the 3D-Future dataset. Visual results on the 3D-Future dataset, showing that our ViewMask-1-to-3 produces more consistent and realistic novel views compared with recent methods.
Datasets
Our model was trained on a subset of Objaverse [deitke2023objaverse] and HSSD [khanna2023hssd]. Following the data split provided by Objaverse++ [lin2025objaversecurated3dobject], we removed samples labeled as Low Quality, to ensure more stable training. To enrich the textual supervision, we adopted the Cap3D [luo2023scalable, luo2024view] annotations, which provide descriptive captions for 3D objects. In total, we used 180K 3D objects, each rendered as an 8-frame RGBA orbit sequence at a resolution of 256×256, with the elevation fixed at 30°. During training, one frame was randomly selected as the reference image for each object, resulting in a multi-view generation dataset.
Metrics
We evaluate the multi-view generation results from two perspectives. First, for the directly generated novel views, we adopt common image-level metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [1284395], and Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], to assess visual fidelity and perceptual quality. Then, we employ InstantMesh [xu2024instantmesh] to reconstruct 3D geometry from the generated views, and evaluate the resulting point clouds using Chamfer Distance (CD) and Intersection over Union (IoU) to quantify geometric consistency and accuracy.
Experimental Settings
[…]
4.1 Image to Multi-view Images
Table 1: Quantitative comparison on multi-view generation over the GSO and 3D-FUTURE datasets. We report PSNR, SSIM, and LPIPS.
Method Architecture Control GSO 3D-FUTURE Rank
[…]
Zero-1-to-3 [liu2023zero] 2D Cont. Diff. Camera Poses 18.8219 0.8294 0.1659 17.0526 0.8163 0.1760 5.3
Zero-1-to-3 XL [liu2023zero, objaverseXL] 2D Cont. Diff. Camera Poses 19.6839 0.8381 0.1518 18.4702 0.8337 0.1485 3.2
ViVid-1-to-3 [kwak2023vivid] 2D Cont. Diff. Camera Poses 19.7978 0.8566 0.1764 18.3241 0.8437 0.1682 3.5
Zero123++ [shi2023zero123++] 2D Cont. Diff. Fixed Views 19.6373 0.8045 0.3550 23.5001 0.8527 0.1529 4.3
Epidiff [huang2024epidiffenhancingmultiviewsynthesis] 2D Cont. Diff. Fixed Views 18.9917 0.8244 0.1667 17.4592 0.8108 0.1785 5.7
MV-Adapter [huang2024mvadapter] 2D Cont. Diff. Fixed Views 19.4673 0.7518 0.1503 19.3375 0.7217 0.4036 6.0
AR-1-to-3 [zhang2025ar] Video Cont. Diff. Fixed Views 13.5084 0.7376 0.3514 13.7724 0.7479 0.3034 8.5
TRELLIS [xiang2024structured] 3D Rectified Flow 3D Asset Render 14.4493 0.8058 0.3223 14.4604 0.7896 0.2990 7.2
[…]
Table 2: Quantitative comparison on 3D reconstruction over the GSO and 3D-FUTURE datasets. We report chamfer distance (CD) value and IOU. $\uparrow$ indicates higher is better, $\downarrow$ indicates lower is better.
Method GSO 3D-FUTURE
[…]
Zero-1-to-3 0.0163 0.5665 0.0113 0.5005
Zero-1-to-3 XL 0.0159 0.5799 0.0106 0.5271
Zero123++ 0.0163 0.5832 0.0314 0.4250
ViVid-1-to-3 0.0163 0.5841 0.0105 0.5246
AR-1-to-3 0.0372 0.4380 0.0148 0.4286
EpiDiff 0.0166 0.5457 0.0121 0.4784
TRELLIS 0.0496 0.4026 0.0178 0.4396
[…]
Baseline
We compare our method against a comprehensive set of state-of-the-art diffusion-based novel view generation approaches, including Zero-1-to-3 [liu2023zero], Zero-1-to-3 XL [liu2023zero, objaverseXL], ViVid-1-to-3 [kwak2023vivid], Zero123++ [shi2023zero123++], MV-Adapter [huang2024mvadapter], EpiDiff [huang2024epidiffenhancingmultiviewsynthesis], AR-1-to-3 [zhang2025ar], and TRELLIS [xiang2024structured].
Qualitative Results
Fig. 4 presents a visual comparison between our method and several state-of-the-art multi-view generation approaches. In the top three rows, ViewMask-1-to-3 preserves accurate chair legs and seat curvature, avoiding the distortions and artifacts seen in other methods. In the bottom three rows, it achieves the best overall consistency, capturing both the recessed details on the desktop and the metal frame. Compared with these baselines, our method produces sharper visual quality, better preserves fine-grained textures, and maintains higher cross-view consistency.
Quantitative Results
To quantitatively evaluate the zero-shot generalization capability of our model, we conduct experiments on two aspects: (1) image-level quality, measured on novel views generated from a single input image, and (2) 3D-level geometric accuracy, evaluated on 3D reconstructions. These experiments are evaluated on the Google Scanned Objects (GSO) dataset [xin2025luminadimooomnidiffusionlarge] and the 3D-FUTURE dataset [fu20203dfuture3dfurnitureshape].
The input views are selected with controlled offsets in azimuth and elevation relative to the canonical frontal view of each 3D object. To account for inconsistent pose assumptions across baselines, we render ground-truth multi-view images using each model’s default viewing configuration. For models that support arbitrary-angle rendering, such as Zero-1-to-3 and Zero-1-to-3 XL, the views are rendered at the same angles as those used by ViewMask-1-to-3 to ensure a fair comparison. We then evaluate PSNR, SSIM, and LPIPS following each model’s respective settings, and report the results in Tab. 1. Overall, our method, ViewMask-1-to-3, achieves the best combined performance across both datasets and attains the highest rank (1.7), demonstrating its superiority in multi-view generation.
To evaluate geometric consistency, we reconstruct 3D shapes from the generated multi-view images for each model. We then compute CD and IoU between the reconstructed shapes and the ground-truth meshes. The quantitative results are reported in Tab. 2, demonstrating that our method achieves lower CD and higher IoU, indicating more accurate 3D reconstruction and better cross-view consistency.
The performance of Zero123++ on the 3D-FUTURE dataset is lower than expected mainly due to its default generation of opaque images with a gray background (corresponding to the zero value in the Stable Diffusion VAE). Removing this background using rembg may inadvertently eliminate fine-grained structures in 3D-FUTURE objects, thereby degrading reconstruction quality. For TRELLIS, some generated 3D assets include an additional bottom platform, which reduces the effective object scale. The necessary post-processing for platform removal and alignment with the ground-truth (GT) geometry introduces additional geometric errors, leading to inferior quantitative results.
4.2 Text to Multi-view Images
Figure 5: Qualitative results on T2MV. Examples of multi-view object generation from text prompts using our method, achieved directly without any intermediate steps.
In the T2MV task, we demonstrate the ability to generate multi-view images of an object directly from text inputs. Models like MV-Adapter [huang2024mvadapter] and TRELLIS [xiang2024structured] support T2MV and I2MV using different checkpoints or architectures. Zero123++ [shi2023zero123++] and ViVid-1-to-3 [kwak2023vivid] generate reference images with a T2I model before extracting the subject for novel views. In contrast, our method directly produces novel views that preserve both the specified viewpoint and the prompt’s semantic content. As shown in Fig. 5, we present examples from the Objaverse dataset that were selected as test samples.
4.3 Ablation study
Figure 6: Qualitative results of different mask scheduling strategies on a subset of the GSO dataset. Compared with linear, quadratic, and fixed schedules, the cosine schedule produces sharper details and more coherent structures in the generated novel views.
Mask Schedule
We investigate the impact of different mask scheduling strategies on our model’s performance, including linear, quadratic, and cosine schedules.
In Tab. 3, our experiments on a subset of the GSO dataset indicate that the cosine schedule yields smoother optimization and superior final results compared with the other strategies, highlighting the importance of gradually adjusting the masking ratio over timesteps. In addition to quantitative evaluation, we provide qualitative visualizations of the generated images under different schedules, as shown in Fig. 6. For the boot, both the linear and quadratic schedules hallucinate an additional zipper on the opposite side, which is clearly inconsistent with reality. For the chair, the quadratic schedule yields incorrect thickness estimation. For the toy, both the linear and quadratic schedules distort the teddy bear’s facial region.
Table 3: Ablation study on different mask scheduling strategies. The metrics (e.g., PSNR, SSIM, LPIPS) are reported for each schedule.
Schedule PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Linear 17.9882 0.8435 0.1882
Quadratic 17.9785 0.8431 0.1865
Cosine 18.0974 0.8437 0.1834
4.4 Task Extension
Figure 7: Task Extension. Left: basic multimodal understanding (MMU), e.g., captioning. Middle: text-to-image (T2I) generation without fine-tuning. Right: training-free 2D turn-style and completion preserving the background.
Unifed Multimodal Understanding and Generation
In addition to the aforementioned image-to-multiview images(I2MV) and text-to-multi-view images(T2MV) capabilities, the Stage-1 pretraining endows our model with fundamental multimodal understanding (MMU) and text-to-image (T2I) generation abilities. As shown in Fig. 7, the model can perform basic understanding and generation tasks without any additional fine-tuning. Building upon these foundational abilities, our model can naturally extend toward a unified multimodal understanding and generation framework, enabling effortless adaptation to a wide range of new multimodal tasks.
Training-Free 2D Turn-Style
Inspired by Adobe’s Project Turn Style, which enables re-angling and rotating 2D illustrations as if they were 3D objects, we extend this capability in a training-free manner based on our ViewMask-1-to-3 framework. Benefiting from the inherent properties of discrete diffusion, our model can rotate a specified object by a given angle and perform image completion while preserving the original background, all without additional fine-tuning, as these abilities are implicitly embedded during diffusion training. These cases demonstrate that our framework not only extends toward a unified diffusion architecture, but also exhibits strong flexibility and generalization across diverse generation and reasoning tasks.
5 Conclusion
In this work, we introduce ViewMask-1-to-3, a discrete diffusion framework for multi-view image generation. Experiments on GSO and 3D-FUTURE demonstrate strong performance, highlighting discrete diffusion as an effective and flexible alternative to traditional continuous diffusion and 3D-aware approaches. Our results also suggest the model’s potential for general multimodal understanding (MMU), text-to-image (T2I) generation, text-to-multi-view (T2MV) synthesis, and training-free 2D turn-style image manipulation, pointing to a possible path toward more unified and scalable multi-view generation.
Despite its effectiveness, limitations remain, such as restricted elevation angles, a small number of viewpoints, and differences in lighting during turn-style generation. Future work will explore increasing viewpoint diversity, improving resolution, and better handling lighting variations to further advance unified and scalable multi-view generation.
[…]
Supplementary Material
Appendix A Additional Experimental Details
A.1 Baseline Setting
Figure 8: Perspective vs. Orthographic Projection.
As shown in Fig. 8, most of the baselines we select—such as Zero-1-to-3 [liu2023zero], Zero-1-to-3 XL [liu2023zero, objaverseXL], ViVid-1-to-3 [kwak2023vivid], Zero123++ [shi2023zero123++], EpiDiff [huang2024epidiffenhancingmultiviewsynthesis], AR-1-to-3 [zhang2025ar], TRELLIS [xiang2024structured], and our own method ViewMask-1-to-3 are trained on data rendered with perspective projection. Among all these methods, only MV-Adapter [huang2024mvadapter] is trained using data rendered with orthographic projection.
A.2 Ground-Truth Rendering
For Zero123++ and AR-1-to-3, the ground-truth images are rendered from 3D assets using their default six viewpoint configuration. MV-Adapter, follows its original training protocol and renders four orthographic views consistent with its orthographic camera model. For the remaining baselines that support arbitrary multi-view rendering, we use the same viewpoint and camera parameters as ours to ensure fair comparison.
For models such as Zero123++ that by default generate images with a gray background, we render the corresponding ground-truth views using the same background color when reporting PSNR, SSIM, and LPIPS. Although some implementations offer background removal via rembg, this often removes thin or delicate parts of the object as well, which in turn compromises the reliability of the baseline’s quantitative results.
Appendix B Additional Qualitative Comparison
B.1 Image to Multi-view Images
We provide additional I2MV comparison results, as shown in Fig. 10. Our model demonstrates more accurate control over fine-grained details.
Figure 9: Qualitative results on T2MV. Examples of multi-view object generation from text prompts using MV-Adapter, Trellis, ViewMask-1-to-3.
Figure 10: Qualitative comparison on the 3D-Future dataset. Visual results on the 3D-Future dataset, showing that our ViewMask-1-to-3 produces more consistent and realistic novel views compared with recent methods.
B.2 Text to Multi-view Images
To further illustrate the qualitative performance differences, we visualize and compare the results produced by MV-Adapter [huang2024mvadapter], TRELLIS [xiang2024structured], and our method, as shown in Fig. 9. In contrast to MV-Adapter, whose orthographic projection setup limits the amount of recoverable geometric detail, our method produces richer visual details across views. Compared with TRELLIS, which represents one of the strongest baselines, our approach achieves comparable visual quality while requiring only a single inference pass and using the same unified checkpoint across all tasks.
B.3 Task Extension
Unified Multimodal Understanding and Generation
During inference, ViewMask employs task-specific input templates that leverage discrete tokens for both vision and language modalities. For multimodal understanding (MMU) tasks, the input sequence is structudarkred as:
[…]
where the text prompt and visual content are presented sequentially to guide the model’s comprehension and response generation. The red-highlighted tokens correspond to the fully masked answer region, representing the content the model must produce.
For text-to-image (T2I) task, the sequence follows:
$\text{[T2I] [SOT] }\textit{caption}\text{ [EOT] [SOI] }{\color[rgb]{0.62890625,0,0}\textit{image tokens}}\text{ [EOI]},$
where the model autoregressively generates image tokens after the caption prompt.
Training-Free 2D Turn-Style
As shown in Fig. 7, ViewMask-1-to-3 implements a similar training-free capability, allowing 2D illustrations to be re-angled as if they were 3D. Our pipeline performs two sequential inference passes to achieve multi-view image generation while preserving the original scene context. In the first pass, the model receives the object obtained from the original image and generates its appearance across multiple viewpoints. In the second pass, the object is masked and provided to the model again to reconstruct the masked regions, restoring the original background while integrating the previously generated object views. By fusing the outputs of both stages, the final results yield consistent multi-view objects seamlessly incorporated into the original scene, all achieved without additional fine-tuning.